Circular No. 6755

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1997 T3

Further precise CCD positions have been reported as follows:

1997 UT α_{2000} δ_{2000} m_{2} Observer
Oct. 6.99668 1^{h}14^{m}23.^{s}30 +2^{o}59′08′′.2 19.0 Tichý
7.00264 1 14 23.13 +2 59 07.4  _{”}
7.00669 1 14 23.00 +2 59 06.6  _{”}
7.28978 1 14 14.58 +2 58 25.2 18.8 Offutt
7.34699 1 14 12.83 +2 58 16.2 18.5  _{”}
7.38441 1 14 11.71 +2 58 10.0 18.5  _{”}
7.87075 1 13 57.21 +2 56 57.2  Pravec
7.90740 1 13 56.08 +2 56 52.1  _{”}
7.93979 1 13 55.10 +2 56 46.9 19.0  _{”}

M. Tichý and Z. Moravec (Kleť). 0.57-m f/5.2 reflector. No evident coma or tail. Comparison with the Bochum images suggests the tail was too faint to show, even on coaddition of the Kleť images.

W. Offutt (Cloudcroft). 0.6-m f/7 Ritchey-Chrétien. No coma or tail immediately apparent. Subsequent careful inspection consistently revealed a possible trace of tail, up to 17′′ in p.a. 247^{o}.

P. Pravec (Ondřejov). 0.65-m f/3.6 reflector. No coma, but very faint tail on individual 3-min clear-filter exposures, the coaddition of three images confirming a tail 25′′ long in p.a. 240^{o}.

The general elliptical orbital elements below are from the 16 observations above and on IAUC 6754. The observations can also be satisfied by a low-eccentricity Jupiter-Trojan orbit, though not, it seems, by a parabola.

T = 1997 Nov. 1.809 TT  ω = 311.^{o}119
e = 0.49983  Ω = 70.171 }2000.0
q = 3.76725 AU  i = 4.162

a = 7.53198 AU n^{o} = 0.047680 P = 20.67 years

1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Sept. 29 1^{h}18.^{m}24 + 3^{o}19.^{′}4 2.794 3.773 165.^{o}9 3.^{o}7 19.0
Oct. 4 1 15.87 + 3 06.8 2.780 3.772 170.9 2.4 18.9
9 1 13.39 + 2 54.2 2.774 3.770 175.0 1.3 18.8
14 1 10.86 + 2 42.1 2.775 3.769 174.4 1.5 18.8
19 1 08.37 + 2 30.9 2.783 3.768 170.0 2.6 18.9
24 1 05.96 + 2 20.9 2.799 3.768 164.8 4.0 19.0
29 1 03.70 + 2 12.5 2.821 3.767 159.5 5.3 19.1
1997 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden